OpenStudy (anonymous):

Suppose that 2400 is invested at an interest rate of 2 %per year, compounded continuously. After how many years will the initial investment be doubled?

OpenStudy (anonymous):

A(t) = P e^(rt) A = total money after t years P = initial investment r = rate t = time A = 2x P = x r = 0.02 t = ? 2x = x e^(0.02t) 2 = e^(0.02t) ln 2 = 0.02t \[\frac{\ln2}{0.02}= t\]